# Clarifying Kasner dynamics inside anisotropic black hole with vector hair

### v2

@@ 1 Introduction
The interior dynamics for the black holes with vector hair were considered, which is more challenging than the scalar case due to the spatial anisotropy. Before reaching the singularity, several intermediate regimes develop, including the Einstein-Rosen bridge contracting towards the singularity, the oscillations of vector condensate and the anisotropy of spatial geometry, and a sequence of Kasner epochs connected by bounces. The mechanism for the former two behaviors were clearly explained Cai:2021obq . Nevertheless, the alternation of Kasner epochs was only observed numerically Cai:2021obq ; Sword:2022oyg . Although it was described geometrically as collisions between electric “walls” Henneaux:2022ijt (see also DeClerck:2023fax ), a comprehensive understanding of the dynamics has not yet to be established. In particular, the empirical relations for the alternation of Kasner epochs were reported in Cai:2021obq summarized by a large number of numerical examples. Near some specific temperature values, there develop numerous oscillations extending deep towards the black hole singularity with the amplitude of oscillations increasing Sword:2022oyg . The interior could yield further interesting behaviors upon exploring the parameter space. This paper aims to provide insights into the internal structure of anisotropic black holes with vector hair and to give a clear interpretation on the numerical results reported in Cai:2021obq ; Sword:2022oyg . More precisely, we will consider the charged vector case known as holographic p-wave superconductor Cai:2013pda ; Cai:2013aca . Based on the method developed in our previous work Cai:2023igv , we will classify three different types of Kasner alternations analytically, including the Kasner transition, Kasner inversion and Kasner reflection. Our analytical results will be corroborated by numerical solutions to the full equations of motion.
@@ 2.2 Proof of No Inner Horizon
There is a scaling symmetry associated with $\mathcal{S}_{eff}$.
$z\to\lambda z\,,\quad h\to h\,,\quad\chi\to\chi-2d\ln{\lambda}\,,\quad\zeta\to\zeta+\frac{\ln{\lambda}}{d-2}\,,\quad A_{t}\to\lambda^{d-1}A_{t}\,,\quad\rho_{x}\to\rho_{x}\,,$ (9)
@@ 3.1 Kasner Solution
It is easy to show that in the Kasner universe (17) one metric component grows while others shrink for $d=3$ case, while at least one metric component grows for higher dimensions. Nevertheless, the overall spatial volume density collapses as approaching the singularity.
@@ 3.3 Kasner Transition and Reflection
since in this case $A_{t}$, $h$ and $\rho_{x}$ are finite.
Therefore, when $\beta<1/(d-2)$, we are able to estimate the order of $\rho_{x}^{\prime}$:
@@
Therefore, both terms are negligible. The spacetime will stay in a stable Kasner epoch (15) if $h^{\prime}$ is also integrable.
@@
with $c_{2}$ and $z_{R}$ integration constants. Similar to the Kasner transition, $\widetilde{\beta}$ becomes a constant far away from $z_{R}$, corresponding to a Kasner epoch described by (15). The alternation of the two Kasner epochs is triggered near $z_{R}$.
It is easy to obtain the relation between the two adjacent Kasner epochs:
@@ 3.4 Laws of Kasner Transformation
$\begin{cases}\mathrm{Kasner\ transition}:\beta+\beta_{T}=\frac{2}{d-2},\quad\beta>\frac{1}{d-2}\,,\\ \mathrm{Kasner\ inversion}:\quad\beta\,\beta_{I}=1,\quad-1<\beta<1\,,\\ \mathrm{Kasner\ reflection}:\beta+\beta_{R}=-2,\quad\beta<-1\,.\\ \end{cases}$ (50)
@@
As clearly from (18), all Kasner exponents of a Kasner universe are fully characterized by the parameter $\beta$. One can see from (50) that our laws cover all parameter space of $\beta$, suggesting that we have clarified all possible kinds of Kasner transformation for the anisotropic black holes with vector hair (2).
@@
Interestingly, we find that while the law of Kasner transition depends on the dimension of spacetime, the laws of Kasner inversion and Kasner reflection are dimension-independent. In particular, when $d=3$, there is no overlapping region between the Kasner inversion and transition, for which the transformation laws (50) will be valid for any Kasner epoch, except for the critical case with $|\beta|=1$. However, when $d>3$, there may be competing behaviors between the Kasner inversion and transition in the overlapping region $\frac{1}{d-2}<\beta<1$, which may beyond our laws of (50). So far, we have not been able to give an analytical description of this overlapping regime. We shall give more details in Section 3.5. Nevertheless, the transformation laws (50) suggest the chaotic evolution characterized by the random change of the amplitude of the Kasner exponent at late interior times. In particular, independent of the spacetime dimension, there exist always the fixed point $\beta=-1$ approaching which numerous Kasner epochs will be triggered with the Kasner exponent $\beta$ been amplified continuously towards the singularity (see Section 4.2).
@@ 3.5 Numerical Check
with the dots standing for the higher order terms in the expansion of $z$. For definiteness, we limit ourselves to the case for which the vector hair develops spontaneously, i.e. $\rho_{s}=0$. It corresponds to the holographic p-wave superconductor Cai:2013pda ; Cai:2013aca where $\rho_{v}$ is the condensate and $\mu$ is the chemical potential. Without loss of generality, we shall fix the chemical potential to be $\mu=1$. Then we are left with only one-dimensional parameter space of solutions to explore, which can be labelled by the temperature. More precisely, below a critical temperature $T_{c}$, the vector condensate develops smoothly, breaking the U(1) symmetry as well as rotational symmetry spontaneously, see Cai:2013aca ; Cai:2014ija for more details.
@@ 3.5.1 Case with $d=3$
Figure 2: The black hole interior demonstrating function $z\zeta^{\prime}$ at $T\approx 0.982T_{c}$ (left) and $T\approx 0.975T_{c}$ (right). The Kasner exponent $\beta$ at each Kasner epoch is marked, from which one can verify the transformation laws (50) directly. The shaded region denotes the overlapping regime for the Kasner transition and Kasner inversion. The fixed point with $\beta=-1$ is denoted by the green-dashed line. We have chosen $d=5$, $q=\sqrt{10}$ and $L=1$.
@@ 3.5.2 Case for $d>3$
For $d>3$, there is an overlap between the Kasner transition and inversion processes when $11/(d-2)<\beta<$ 1. For the sake of specificity, we consider the case with $d=5$ for which the transformation law reads
$\begin{cases}\mathrm{Kasner\ transition}:\beta+\beta_{T}=\frac{2}{3},\quad\beta>\frac{1}{3}\,,\\ \mathrm{Kasner\ inversion}:\quad\beta\,\beta_{I}=1,\quad-1<\beta<1\,,\\ \mathrm{Kasner\ reflection}:\beta+\beta_{R}=-2,\quad\beta<-1\,.\\ \end{cases}$ (59)
The parameter space has a overlap when $11/3<\beta<$ 1, corresponding to the shaded region between the blue dotted and red dotted lines in Figure 2. In this overlapping regime, the complex competition between the Kasner transition and inversion could occur, for which we cannot give a clear analytical description. In the left panel, the third Kasner epoch with $\beta_{3}=0.8955$ lies within the overlapping regime from a Kasner reflection. Then there could be a Kasner transition, a Kasner inversion or beyond both. Our numerical computation shows the develop of the next Kasner epoch with $\beta_{4}=-0.2288$. One can check that $\beta_{3}\beta_{4}=-0.2049$ and $\beta_{3}+\beta_{4}=0.6667$. Therefore, the alternation is well described by the Kasner transition. In the right panel, one can see the seventh Kasner epoch with its component $\beta_{7}=0.4431$ that is within the overlapping regime. The next Kasner epoch has the exponent $\beta_{8}=2.2613$. One then finds that $\beta_{7}\beta_{8}=1.0020$ and $\beta_{3}+\beta_{4}=2.7044$. Thus, the alternation between the above two Kasner epochs is well captured by the Kasner inversion.
Based on a large number of numerical examples, we observe that, for a Kasner epoch within the range $\frac{1}{d-2}<\beta<1$, the larger the value of $\beta$, the more likely the Kasner transition is to occur. Otherwise, the smaller the $\beta$, the more likely the Kasner inversion is to develop. It remains uncertain whether there exists an exact boundary that determines which transformation law is dominant. Moreover, it might develop an epoch beyond the Kasner universe in the overlapping regime at some temperatures.
In contrast, outside the overlapping region, the alternation of adjacent Kasner epochs are well predicted by our laws (50), which is manifest in Figure 2. Our analysis suggests that there exhibits a never-ending oscillatory behavior towards the spacelike singularity.
@@ 4.1 Alternation of Kasner Epochs
where we have used the asymptotic solution of $\zeta\sim\beta\ln z$.
For the 4-dimensional case reported in Cai:2021obq , one can then easily verify that the two Kasner transformation laws they found are:
@@ 5 Conclusion and Discussion
Although we have managed to explain the key dynamics of Kasner epochs at later interior times, the parameter space of Kasner inversion and the one of Kasner transition have an overlap when $\frac{1}{d-2}<\beta<1$. Within this range, the competition between the Kasner inversion and Kasner transition appears, see Figure 2. Whether there exists an exact boundary that determines which transformation law is dominant is unclear. In the present study, we have limited to the massless vector field. However, our analysis can be generalized to massive vector field. While the mass term was argued to be negligible in the far interior dynamics for neutral vector field DeClerck:2023fax , it appears that the mass term $m^{2}\rho^{\mu}\rho_{\mu}^{\dagger}$ becomes important once the exponent $n_{u}>2(d-1)$ Cai:2021obq , or equivalently, $\frac{2\beta}{1+\beta}>d-1$, see (63). Therefore, when the exponent $\beta$ of a Kasner epoch falls into the range $-\frac{d-1}{d-3}<\beta<-1$, the contribution from the mass term could affect the Kasner dynamics, in particular, for the Kasner reflection.  [^4] Meanwhile, the mass term seems to play a key role using cosmological billiards formulated as a hyperbolic billiards problem Henneaux:2022ijt . It will be interesting to understand the role of massive term and other general interactions in the interior dynamics. There have been some studies on the internal structure of black holes with rotation Gao:2023rqc and higher-derivative corrections Bueno:2024fzg ; Devecioglu:2023hmn . Extending our approach to these cases would be interesting. Moreover, it will be desirable to understand the holographic meaning of the black hole interior and to find a way to probe the interior dynamics Caceres:2023zft ; Mansoori:2021wxf ; Ceplak:2024bja ; Arean:2024pzo . We hope to return to these issues in future work.

### v3

@@ 1 Introduction
The interior dynamics for the black holes with vector hair were considered, which is more challenging than the scalar case due to the spatial anisotropy. Before reaching the singularity, several intermediate regimes develop, including the Einstein-Rosen bridge contracting towards the singularity, the oscillations of vector condensate and the anisotropy of spatial geometry, and a sequence of Kasner epochs connected by bounces. The mechanism for the former two behaviors were clearly explained Cai:2021obq . Nevertheless, the alternation of Kasner epochs was only observed numerically Cai:2021obq ; Sword:2022oyg . Although it was described geometrically as collisions between p-form “walls” Henneaux:2022ijt (see also DeClerck:2023fax ), a comprehensive understanding of the dynamics has not yet to be established. In particular, the empirical relations for the alternation of Kasner epochs were reported in Cai:2021obq summarized by a large number of numerical examples. Near some specific temperature values, there develop numerous oscillations extending deep towards the black hole singularity with the amplitude of oscillations increasing Sword:2022oyg . The interior could yield further interesting behaviors upon exploring the parameter space. This paper aims to provide insights into the internal structure of anisotropic black holes with vector hair and to give a clear interpretation on the numerical results reported in Cai:2021obq ; Sword:2022oyg . More precisely, we will consider the charged vector case known as holographic p-wave superconductor Cai:2013pda ; Cai:2013aca . Based on the method developed in our previous work Cai:2023igv , we will classify three different types of Kasner alternations analytically, including the Kasner transition, Kasner inversion and Kasner reflection in general spacetime dimensions. This provides a different way from the cosmological billiard to understand the black hole interior dynamics. In $(3+1)$ dimensional spacetime, the corresponding transformation rules for the Kasner exponent agree with the known formula obtained via the cosmological billiard, confirming the correctness of our analysis. We will present numerical checks by solving the full equations of motion. We manage to provide a clear explanation of the alternation of Kasner epochs found numerically in the literature Cai:2021obq ; Sword:2022oyg .
@@ 2.2 Proof of No Inner Horizon
$z\to\lambda z\,,\quad h\to h\,,\quad\chi\to\chi-2d\ln{\lambda}\,,$ (9)
$\zeta\to\zeta+\frac{\ln{\lambda}}{d-2}\,,\quad A_{t}\to\lambda^{d-1}A_{t}\,,\quad\rho_{x}\to\rho_{x}\,,$
@@ 3.1 Kasner Solution
It is easy to show that in the Kasner universe (17) one metric component grows while others shrink for $d=3$ case, while at least one metric component grows for higher dimensions. We will see this feature more clearly in Sec. 3. Nevertheless, the overall spatial volume density collapses as approaching the singularity.
@@ 3.3 Kasner Transition and Reflection
since in this case $A_{t}$, $h$ and $\rho_{x}$ are finite. Therefore, when $\beta<1/(d-2)$, we are able to estimate the order of $\rho_{x}^{\prime}$:
@@
One should pay more attention when $11/(d-2)<\beta<$ 1 for which there develops an overlapping region between the Kasner inversion and transition, and there may be competing behaviors between the Kasner inversion and transition in the overlapping region. To solve this issue, we come back to (3) and estimate the contribution of each term. In particular, we need to compare the following two terms: $(z\rho_{x}^{\prime})^{2}\mathrm{e}^{-2(d-2)\zeta}$ and $h^{\prime}\zeta^{\prime}/h$. We then find that there is a critical Kasner exponent
$\beta_{c}=-1+\sqrt{\frac{2(d-1)}{d-2}}\,,$ (44)
for which
$\mathcal{O}\left((z\rho_{x}^{\prime})^{2}\mathrm{e}^{-2(d-2)\zeta}\right)=\mathcal{O}\left(\frac{\mathrm{e}^{2(d-2)\zeta}}{z^{4}}\right)=\mathcal{O}\left(\frac{h^{\prime}\zeta^{\prime}}{h}\right)\,.$ (45)
Therefore, when $\beta>\beta_{c}$, the term $(z\rho_{x}^{\prime})^{2}\mathrm{e}^{-2(d-2)\zeta}$ regarding the Kasner transition dominates the dynamics. In contrast, when $\beta<\beta_{c}$ the Kasner alternation is governed by the Kasner inversion.
@@
Therefore, both terms are negligible. Hence, combined with the above analysis in subsection 3.2, Kasner inverison will dominate the transformation process in this $\beta$ region.
@@
with $c_{2}$ and $z_{R}$ integration constants. Similar to the Kasner transition, $\widetilde{\beta}$ becomes a constant far away from $z_{R}$, corresponding to a Kasner epoch described by (15). The alternation of the two Kasner epochs is triggered near $z_{R}$. It is easy to obtain the relation between the two adjacent Kasner epochs:
@@ 3.4 Laws of Kasner Transformation
$\begin{cases}\mathrm{Kasner\ transition}:\beta+\beta_{T}=\frac{2}{d-2},\quad\beta>\beta_{c}\,,\\ \mathrm{Kasner\ inversion}:\quad\beta\,\beta_{I}=1,\quad-1<\beta<\beta_{c}\,,\\ \mathrm{Kasner\ reflection}:\beta+\beta_{R}=-2,\quad\beta<-1\,,\\ \end{cases}$ (52)
@@
where $\beta_{c}=-1+\sqrt{2(d-1)/(d-2)}$. As clearly from (18), all Kasner exponents of a Kasner universe are fully characterized by the parameter $\beta$. One can see from (52) that our laws cover all parameter space of $\beta$, suggesting that we have clarified all possible kinds of Kasner transformation for the anisotropic black holes with vector hair (2).
@@
Interestingly, we find that while the law of Kasner transition depends on the dimension of spacetime, the laws of Kasner inversion and Kasner reflection are dimension-independent. In particular, when $d=3$, there is a symmetric region between the Kasner reflection and transition, for which the transformation laws (52) are completely same with the situation of pure gravity Lifshitz:1963ps ; Belinsky:1970ew and scalarized case An:2022lvo . The transformation laws (52) suggest the chaotic evolution characterized by the random change of the amplitude of the Kasner exponent at late interior times. In particular, independent of the spacetime dimension, there exist always the fixed point $\beta=-1$ approaching which numerous Kasner epochs will be triggered with the Kasner exponent $\beta$ been amplified continuously towards the singularity (see Section 4.2).
Besides, we find that the Kasner transformation and the minimum negative value of the Kasner exponents is closely correlated. Considering the expressions of Kasner exponents (18), we can obtain
$\mathrm{Min}\{p_{t},p_{x},p_{\sigma}\}=\begin{cases}p_{x}<0,\ \mathrm{Kasner\ transition}:\beta>\beta_{c}\,,\\ p_{t}<0,\ \mathrm{Kasner\ inversion}:-1<\beta<\beta_{c}\,,\\ p_{\sigma}<0,\ \mathrm{Kasner\ reflection}:\beta<-1\,.\\ \end{cases}$ (53)
It suggests that, in different $\beta$ regions, the corresponding minimum and negative Kasner exponent leads to the instability of Kasner space-time, and then the Kasner transformation.
It is interesting to compare our approach to the the cosmological billiard method in terms of an elegant geometric description. The basic elements of the latter is to know the specific locations of the walls (i.e. the boundaries of the billiard table), and the reflection rules about the walls. The corresponding transformation rules for the Kasner exponent have been worked out in general (without symmetry assumptions) in Demaret:1986ys ; DEMARET198527 for gravitational walls and in Damour:2000wm ; Damour:2002fz for p-form walls. In $d=3$, the gravitational and the electromagnetic walls coincide and there is no symmetry wall when the metric is diagonal. Therefore, the billiard is the ideal triangle delimited by the gravitational walls. The collision against a gravitational wall yields as the following Kasner exponents Damour:2002et
$p_{1}^{\prime}=-\frac{p_{1}}{1+2p_{1}},\quad p_{2}^{\prime}=\frac{p_{2}+2p_{1}}{1+2p_{1}},\quad p_{3}^{\prime}=\frac{p_{3}+2p_{1}}{1+2p_{1}}\,,$ (54)
where $\{p_{1},p_{2},p_{3}\}$ denote the Kasner exponents with different spatial directions and $p_{1}$ is the smallest Kasner exponent. Note that $\beta_{c}=1$ for $d=3$. From (53), one has $p_{1}=p_{x}$ when $\beta>1$, for which the formula (54) gives our Kasner transition formula. When $|\beta|<1$, it is $p_{t}$ that is the smallest exponent (i.e. $p_{1}=p_{t}$) and the formula (54) gives then the Kasner inversion formula. Finally, when $\beta<-1$, it is $p_{\sigma}$ that is the smallest exponent (i.e. $p_{1}=p_{\sigma}$), for which the formula (54) reduces to the Kasner reflection. This demonstrates the truth of our formula (52) and suggests that our theory could be described by cosmological billiard.
Regarding to our present theory, to the best of our knowledge, there has no explicit formula as (54) via cosmological billiard when $d>3$ in the literature. Indeed, we find that, when $d>3$, our laws of Kasner transformation (54) can be summarized as
$p_{1}^{\prime}=-\frac{p_{1}}{1+\frac{n_{1}^{2}}{d-2}2p_{1}},\quad p_{2}^{\prime}=\frac{p_{2}+\frac{n_{1}^{2}}{d-2}2p_{1}}{1+\frac{n_{1}^{2}}{d-2}2p_{1}},\quad p_{3}^{\prime}=\frac{p_{3}+\frac{n_{1}^{2}}{d-2}2p_{1}}{1+\frac{n_{1}^{2}}{d-2}2p_{1}}\,,$ (55)
with $p_{1}<p_{2}<p_{3}$. Here $n_{1}$ denotes the coefficient of $p_{1}$ in the Kasner relationship $p_{t}+p_{x}+(d-2)p_{\sigma}=1$. More precisely, $n_{1}=1$ when $p_{1}=p_{t}$ or $p_{1}=p_{x}$, and $n_{1}=d-2$ when $p_{1}=p_{\sigma}$. It is interesting to understand the interior dynamics using cosmological billiards and to check if the surviving various types of walls yield the formula we obtain here. Nevertheless, it is beyond the scope of our present study. While we believe that the cosmological billiards method can also handle the Kasner dynamics in the present model, our approach is easy to generalize to more general interactions.
@@ 3.5 Numerical Check
with the dots standing for the higher order terms in the expansion of $z$. For definiteness, we limit ourselves to the case for which the vector hair develops spontaneously, i.e. $\rho_{s}=0$. It corresponds to the holographic p-wave superconductor Cai:2013pda ; Cai:2013aca where $\rho_{v}$ is the condensate and $\mu$ is the chemical potential. Without loss of generality, we shall fix the chemical potential to be $\mu=1$, which means all physical quantities are measured in unit of the chemical potential. Finally, we are left with only one-dimensional parameter space of solutions to explore, which can be labelled by the temperature (or equivalently $T/\mu$). More precisely, below a critical temperature $T_{c}$, the vector condensate develops smoothly, breaking the U(1) symmetry as well as rotational symmetry spontaneously. For the massless vector of (1), the phase transition from the normal phase for which $\rho_{x}=0$ to the broken phase with non-trivial $\rho_{x}$ can be a second order transition or a first order one, depending on the the value of charge $q$. Please consult Cai:2013aca ; Cai:2014ija for more details, in particulate the numerical strategy for constructing the hairy solutions.
@@ 3.5.2 Case for $d>3$
For the sake of specificity, we consider the case with $d=5$ for which the transformation law reads
$\begin{cases}\mathrm{Kasner\ transition}:\beta+\beta_{T}=\frac{2}{3},\quad\beta>\beta_{c}(5)\,,\\ \mathrm{Kasner\ inversion}:\quad\beta\,\beta_{I}=1,\quad-1<\beta<\beta_{c}(5)\,,\\ \mathrm{Kasner\ reflection}:\beta+\beta_{R}=-2,\quad\beta<-1\,,\\ \end{cases}$ (64)
where $\beta_{c}(5)\sim 0.63$. Our numerical computation shows the develop of the next Kasner epoch with $\beta_{4}=-2.3974$. One can check that $\beta_{4}+\beta_{5}=-1.0011$, so the alternation is well described by the Kasner reflection. The seventh Kasner epoch with its component $\beta_{5}=0.3985$. The next Kasner epoch has the exponent $\beta_{6}=2.5121$. One then finds that $\beta_{5}\beta_{6}=1.0011$, so the alternation between the above two Kasner epochs is well captured by the Kasner inversion. The next Kasner epoch has the exponent $\beta_{7}=-1.8460$. One then finds that $\beta_{6}+\beta_{7}=0.6661\sim 2/3$, so the alternation between the above two Kasner epochs is well captured by the Kasner transition. As manifestly shown in Fig. 2, the numerical results are highly consistent with the transition patterns (52) we obtained. Our analysis suggests that there exhibits a never-ending oscillatory behavior towards the spacelike singularity.
Figure 2: The black hole interior demonstrating function $z\zeta^{\prime}$ at $T\approx 0.982T_{c}$ (left) and $T\approx 0.975T_{c}$ (right). The Kasner exponent $\beta$ at each Kasner epoch is marked, from which one can verify the transformation laws (52) directly. The two fixed points with $\beta={-1,\beta_{c}(5)}$ are denoted by the green-dashed and blue-dashed lines. We have chosen $d=5$, $q=\sqrt{10}$ and $L=1$.
@@ 4.1 Alternation of Kasner Epochs
where we have used the asymptotic solution of $\zeta\sim\beta\ln z$. For the 4-dimensional case reported in Cai:2021obq , one can then easily verify that the two Kasner transformation laws they found are:
@@ 5 Conclusion and Discussion
In the present study, we have limited to the massless vector field. However, our analysis can be generalized to massive vector field. While the mass term was argued to be negligible in the far interior dynamics for neutral vector field DeClerck:2023fax , it appears that the mass term $m^{2}\rho^{\mu}\rho_{\mu}^{\dagger}$ becomes important once the exponent $n_{u}>2(d-1)$ Cai:2021obq , or equivalently, $\frac{2\beta}{1+\beta}>d-1$, see (68). Therefore, when the exponent $\beta$ of a Kasner epoch falls into the range $-\frac{d-1}{d-3}<\beta<-1$, the contribution from the mass term could affect the Kasner dynamics, in particular, for the Kasner reflection.  [^5] Meanwhile, the mass term seems to play a key role using cosmological billiards formulated as a hyperbolic billiards problem Henneaux:2022ijt . It will be interesting to understand the role of massive term and other general interactions in the interior dynamics. There have been some studies on the internal structure of black holes with rotation Gao:2023rqc and higher-derivative corrections Bueno:2024fzg ; Devecioglu:2023hmn . Extending our approach to these cases would be interesting. Moreover, it will be desirable to understand the holographic meaning of the black hole interior and to find a way to probe the interior dynamics Caceres:2023zft ; Mansoori:2021wxf ; Ceplak:2024bja ; Arean:2024pzo . Last but no least, all possible kinds of Kasner transformations can be collapsed into a single formula (55) in general spacetime dimensions. For $d=3$, it is precisely the case due the collision against a gravitational wall using the description in terms of cosmological billiards. It is worth understanding (55) from cosmological billiard for which there could be more than just the gravitational walls in higher dimensions. We hope to return to these issues in future work.